# Robust Deterministic Policy Gradient for Disturbance Attenuation and Its Application to Quadrotor Control*

### v3

@@ IV-B Implementation based on TD3 (RDDPG)
    The corresponding target networks $\pi_{\theta^{\prime}}$, $\mu_{\phi^{\prime}}$ for the two actors
@@ IV-B1 Critic update
$L_{\rm critic}(\psi_{i};B):=$ $\frac{1}{|B|}\sum\limits_{(x,u,w,c,x^{\prime})\in B}{(y-Q_{\psi_{i}}(x,u,w))}^{2},$
$i\in\{1,2\}$
@@
$\tilde{u}$ $=\pi_{\theta^{\prime}}(x^{\prime})+{\rm clip}({\epsilon_{1}},-c,c)\ \ \epsilon_{1}\sim\mathcal{N}(0,\sigma)$
$\tilde{w}$ $=\mu_{\phi^{\prime}}(x^{\prime})+{\rm clip}({\epsilon_{2}},-c,c)\ ,\ \epsilon_{2}\sim\mathcal{N}(0,\sigma)$
where $\epsilon_{1},\epsilon_{2}$ are random noise vectors added in order to smooth out the target values, $\mathcal{N}(0,\sigma)$ implies the Gaussian distribution with zero mean and variance $\sigma$, and the clip function is added in order to guarantee the boundedness of the control inputs. The critic’s online parameters $\phi_{1},\phi_{2}$ are updated by the gradient descent step to minimize the loss
@@ IV-B2 Actor update
IV-B3 Target parameters update
@@ IV-B4 Exploration
where $e_{1},e_{2}\sim\mathcal{N}(0,\sigma)$ are exploration noises [4].
IV-B5 Optimizing $\eta$
As an objective function estimate, we use the target $y$ used above. In particular, we simply assume
${J^{{\pi_{\theta^{\prime}}},{\mu_{\phi^{\prime}}}}}\cong\hat{J}:=\frac{1}{{|B|}}\sum\limits_{(x,u,w,r,x^{\prime})\in B}y$
where the target $y$ is defined as before. Then, we can apply
$J\leftarrow$ $(1-{\alpha_{J}})J+{\alpha_{J}}\hat{J}$
$\eta\leftarrow$ $\eta+\alpha_{\eta}J$
where $J$ is an estimate of the cost function $J^{{\pi_{\theta}},{\mu_{\phi}}}$. The basic idea of the update rule for $\eta$ is as follows: if $J>0$, then increase $\eta>0$ so that $J\leq 0$; if $J<0$, then decrease $\eta>0$ so that $J\leq 0$. In this way, we can find an optimal $\eta$ such that $J\leq 0$. The overall algorithm is summarized in Algorithm 1.
@@ IV-B5 Optimizing $\eta$
7:     for  Time step $k$=0,1,2,…$\tau$-1 do
8:         Select actions $u_{k}=\pi_{\theta}(x_{k})+e_{1}$ and $w_{k}=\mu_{\phi}(x_{k})+e_{2}$, where $e_{1},e_{2}\sim\mathcal{N}(0,\sigma)$ are exploration noises.
9:         Compute the cost $c_{k}:=c(x_{k},u_{k},w_{k})$
10:         Observe the next state $x_{k+1}$
11:         Store the transition tuple $(x_{k},u_{k},w_{k},c_{k},x_{k+1})$ in the replay buffer $D$
12:         Uniformly sample a mini-batch $B$ from the replay buffer $\mathcal{D}$
13:         Update critic network:
@@
14:         Update actor networks by the deterministic policy gradient:
@@
15:         Update $J$ and $\eta$:
$J\leftarrow(1-{\alpha_{J}})J+{\alpha_{J}}\hat{J}$
$\eta\leftarrow\eta+\alpha_{\eta}J$
@@
17:         $\theta_{i}^{\prime}\leftarrow\tau\theta_{i}+(1-\tau)\theta_{i}^{\prime}$ , $\phi_{i}^{\prime}\leftarrow\tau\theta_{i}+(1-\tau)\phi_{i}^{\prime}$
@@
19:end for
V Quadrotor application
V-A Quadrotor dynamics model
@@
where $\mathbf{K}_{D}=\text{diag}(K_{D,x},K_{D,y},K_{D,z})$ is a matrix of drag coefficients. Among the four motors, two opposing motors ($1$th and $3$rd) rotate clockwise, while the others ($2$nd and $4$th) rotate counterclockwise. This causes torque $\tau_{k}$ in the roll, pitch, yaw directions
$\tau_{k}=\begin{bmatrix}\tau_{k,\phi}\\ \tau_{k,\theta}\\ \tau_{k,\psi}\end{bmatrix}$
$=\begin{bmatrix}l^{\prime}k_{F}(\text{RPM}_{k,2}^{2}-\text{RPM}_{k,4}^{2})\\ l^{\prime}k_{F}(-\text{RPM}_{k,1}^{2}+\text{RPM}_{k,3}^{2})\\ k_{M}(\text{RPM}_{k,1}^{2}-\text{RPM}_{k,2}^{2}+\text{RPM}_{k,3}^{2}-\text{RPM}_{k,4}^{2})\\ \end{bmatrix}$
where $l^{\prime}=\cos(\frac{\pi}{4})l$ is the scaled arm length, because the motors of the quadrotor are positioned at a 45-degree angle relative to the $x$ and $y$ axes when the yaw angle is 00, $k_{F}$ is the thrust coefficient, and $k_{M}$ is the torque coefficient. The inertia matrix, denoted by the symbol $I$, represents the inertial properties of the quadrotor’s body.
V-B State space
@@
$x_{k}=[\mathbf{e}_{k}^{\text{pos}},\mathbf{e}_{k}^{\text{vel}},\rho_{k},\omega_{k}]^{T}\in\mathbb{R}^{12}$
where
$\mathbf{e}_{k}^{\text{pos}}$ $=\mathbf{p}_{k}^{\text{target}}-\mathbf{p}_{k}$
@@
$\mathbf{e}_{k}^{\text{vel}}$ $=\mathbf{v}_{k}^{\text{target}}-\mathbf{v}_{k}$
@@
$\rho_{k}$ $=[\phi_{k},\theta_{k},\psi_{k}]^{T}\in\mathbb{R}^{3}$
$\omega_{k}$ $=[\omega_{k,\phi},\omega_{k,\theta},\omega_{k,\psi}]^{T}\in\mathbb{R}^{3}$
where $\mathbf{e}_{k}^{\text{pos}}$ is the position error between the target $\mathbf{p}_{k}^{\text{target}}$ and the quadrotor position $\mathbf{p}_{k}$, $\mathbf{e}_{k}^{\text{vel}}$ is the velocity error between the target $\mathbf{v}_{k}^{\text{target}}$ and the quadrotor $\mathbf{v}_{k}$, $\rho_{k}$ is the Euler angle vector, and $\omega$ represents the angular velocity vector of the quadrotor in its body frame. As in the method used in DQN[13], we used the last four time steps of the state as input to the neural network to help the network better capture dynamics in the environment.
V-C Action space
The action of the quadrotor, $u_{k}$, is the speed of four motors $\text{RPM}_{k,i},i\in\{1,2,3,4\}$, which is continuous value between 00 and $21713.714$, as follows:
@@ V-D Cost function
    $c_{tr}=\begin{cases}\zeta&\text{if }||\mathbf{e}_{k}^{\text{pos}}||_{2}\geq 5\text{ or }\phi_{k},\theta_{k}\geq\frac{\pi}{2}\\ 0&\text{else}\end{cases}$
@@
where $\alpha$, $\beta$, $\epsilon$, $\zeta>0$, and $\lambda<0$ are scaling factors. Since $\alpha$ and $\beta$ are positive, the agent learns a policy that minimizes position and velocity errors. If the quadrotor’s angles become too large, it becomes more susceptible to disturbances and may flip easily. Therefore, minimizing the angles cost $c_{\rho}$ is crucial for stability. Additionally, the living cost $c_{l}$ is a negative reward that encourages the quadrotor to maintain stability. The penalty cost $c_{tr}$ is a large positive value applied when the quadrotor fails to track the target—specifically, when the distance between the target and the quadrotor exceeds $5$ meters, or when the quadrotor flips with roll or pitch angles exceeding $\frac{\pi}{2}$. The agent learns a policy to avoid receiving this penalty, which can accelerate the initial learning process.
VI Experiments and results
VI-A Experiments setups
Figure 1: Trajectories for evaluation.
For experiments, we used gym-pybullet-drones environment [17], which is an open-source quadrotor simulator, built using Python and the Bullet Physics engine. This environment provides a modular and precise physics implementation, supporting both low-level and high-level controls. Crazyflie $2.0$ quadrotor is modeled in the simulator, which is used for our experiments. In experiment, we compare RDDPG with DDPG[5], TD3[4], SAC[2], and PPO[3]. These methods utilize a single agent aiming to minimize the cost function $\tilde{c}$. To enhance their robustness against disturbances, random external disturbances were applied during the learning process. From our own experiences, learning under severe disturbances with the four existing methods failed to yield policies that demonstrated reasonable performance. Therefore, we applied only minor disturbances, limited to ±$\pm 5$ [m/s] when applying the four methods.
@@ VI-A Experiments setups
For evaluation, we generate three trajectories as shown in Fig. 1. During evaluation, we introduce random wind disturbances in the $x$ and $y$ directions of the inertia frame. We consider four scenarios where the maximum wind velocity is limited to ±$\pm 5$, ±$\pm 10$, ±$\pm 15$, and ±$\pm 20$ [m/s], which are referred to as the ±$\pm 5$, ±$\pm 10$, ±$\pm 15$, and ±$\pm 20$ wind speed scenarios in the tables. In each scenario, the initial wind speed is randomly set within a given range and updated every $10$ time steps with normally distributed noise. Each algorithm was tested over 500 episodes and evaluated using three metrics: the mean and standard deviation of the total cost per episode, the average $H_{\infty}$ norm, empirically evaluated as $\frac{\tilde{c}(x_{k},u_{k},w_{k})}{w_{k}^{T}w_{k}}$, and the mean and standard deviation of the position error between the quadrotor and the target. All metrics were averaged over 500 episodes.
TABLE I: The mean and standard deviation of the cost
Bold indicates the best performance algorithm Algorithm Trajectory (1) (2) (3) ±$\pm 5$ wind speed scenario RDDPG $\mathbf{-9.20\pm 0.03}$ $\mathbf{-9.11\pm 0.02}$ $\mathbf{-8.94\pm 0.03}$ DDPG $-7.63\pm 0.09$ $-7.67\pm 0.10$ $-7.49\pm 0.12$ TD3 $-4.25\pm 2.47$ $-7.50\pm 1.65$ $-6.00\pm 2.53$ SAC $-7.42\pm 0.98$ $-8.68\pm 0.26$ $-8.34\pm 0.30$ PPO $19.26\pm 12.01$ $7.72\pm 8.37$ $11.03\pm 9.54$$\pm 10$ wind speed scenario RDDPG $\mathbf{-9.00\pm 0.12}$ $\mathbf{-8.96\pm 0.16}$ $\mathbf{-8.73\pm 0.21}$ DDPG $-7.41\pm 0.24$ $-7.41\pm 0.28$ $-7.20\pm 0.43$ TD3 $0.45\pm 5.95$ $-2.20\pm 5.00$ $-0.94\pm 5.00$ SAC $-7.16\pm 1.05$ $-7.94\pm 0.71$ $-7.54\pm 0.87$ PPO $22.89\pm 15.31$ $22.22\pm 15.42$ $22.71\pm 15.64$$\pm 15$ wind speed scenario RDDPG $\mathbf{-8.49\pm 0.52}$ $\mathbf{-8.36\pm 0.71}$ $\mathbf{-8.17\pm 0.62}$ DDPG $-5.95\pm 2.03$ $-5.98\pm 1.90$ $-5.76\pm 1.78$ TD3 $3.02\pm 6.49$ $1.90\pm 6.28$ $2.64\pm 5.45$ SAC $-2.86\pm 4.55$ $-3.70\pm 4.03$ $-3.32\pm 4.13$ PPO $23.63\pm 17.10$ $23.45\pm 16.95$ $23.27\pm 15.77$$\pm 20$ wind speed scenario RDDPG $\mathbf{-5.04\pm 3.90}$ $\mathbf{-4.77\pm 4.07}$ $\mathbf{-4.46\pm 4.03}$ DDPG $0.18\pm 5.56$ $-0.04\pm 5.76$ $-0.51\pm 5.28$ TD3 $3.76\pm 5.32$ $3.27\pm 4.88$ $3.66\pm 4.77$ SAC $0.18\pm 4.64$ $0.49\pm 4.96$ $0.61\pm 4.98$ PPO $20.81\pm 15.58$ $21.15\pm 15.12$ $21.32\pm 16.77$
Figure 2: The cost curve of RDDPG during training. This shows that RDDPG success to minimize the cost as training progresses.
VI-B Results
Fig. 2 shows the sum of costs during the training steps, demonstrating that RDDPG effectively minimizes the cost as training progresses
Table I presents the mean and standard deviation of the total cost sum per episode. Across all scenarios, RDDPG achieved the lowest total cost sum, which demonstrates superior performance compared to other algorithms. Notably, TD3, the baseline algorithm for RDDPG, exhibited significant performance degradation as wind conditions intensified. In contrast, RDDPG maintained a lower cost sum and greater stability, indicating that RDPG enhances the robustness of the DPG algorithm.
Table II describes the mean of $H_{\infty}$ norm at time step $k$ in all wind speed scenarios. As the wind speed decreases, the magnitude of disturbances also diminishes. Since the $H_{\infty}$ norm is inversely proportional to the disturbance magnitude, scenarios with weaker wind conditions tend to exhibit higher $H_{\infty}$ norm values. RDDPG exhibits a smaller $H_{\infty}$ norm than the other algorithms; this implies that RDDPG effectively minimizes the cost in disturbance scenarios
Table III presents the mean and standard deviation of the positional errors between the target and the quadrotor at time step $k$. DDPG performs best in low wind speed scenarios, RDDPG achieves lower errors as wind speed increases. Although DDPG exhibits smaller distance errors, its higher total cost in Table  I suggests that the quadrotor experiences large angular deviations and significant oscillations. These oscillations make the DDPG controller more vulnerable to strong winds, ultimately causing instability in high-wind scenarios.
TABLE II: Empirically evaluated $H_{\infty}$ norm.
Bold indicates the best performance algorithm Algorithm Trajectory (1) (2) (3) ±$\pm 5$ wind speed scenario RDDPG $\mathbf{247.14}$ $\mathbf{284.04}$ $\mathbf{299.59}$ DDPG $7874.49$ $570.36$ $999.50$ TD3 $1421.04$ $573.83$ $820.78$ SAC $708.79$ $749.96$ $593.65$ PPO $8360.81$ $4086.99$ $7548.96$$\pm 10$ wind speed scenario RDDPG $\mathbf{74.17}$ $\mathbf{112.76}$ $\mathbf{84.36}$ DDPG $185.82$ $282.43$ $202.59$ TD3 $422.65$ $331.02$ $351.93$ SAC $209.82$ $144.25$ $123.55$ PPO $2314.39$ $4058.31$ $2308.93$$\pm 15$ wind speed scenario RDDPG $\mathbf{59.29}$ $\mathbf{44.00}$ $\mathbf{41.09}$ DDPG $75.45$ $87.31$ $82.17$ TD3 $216.96$ $147.81$ $216.73$ SAC $103.45$ $75.54$ $89.90$ PPO $6286.04$ $1281.48$ $2862.71$$\pm 20$ wind speed scenario RDDPG $\mathbf{140.33}$ $\mathbf{30.06}$ $\mathbf{53.50}$ DDPG $76.90$ $83.34$ $79.91$ TD3 $284.40$ $205.07$ $176.43$ SAC $88.42$ $73.83$ $78.32$ PPO $1278.23$ $885.98$ $1576.79$
TABLE III: The mean and standard deviation of the position error $||\mathbf{e}_{k}^{pos}||_{2}$ between the target and the quadrotor
Bold indicates the best performance algorithm Algorithm Trajectory (1) (2) (3) ±$\pm 5$ wind speed scenario RDDPG $0.17\pm 0.02$ $0.16\pm 0.02$ $0.17\pm 0.03$ DDPG $\mathbf{0.13\pm 0.05}$ $\mathbf{0.12\pm 0.05}$ $\mathbf{0.12\pm 0.07}$ TD3 $0.35\pm 0.28$ $0.21\pm 0.11$ $0.26\pm 0.17$ SAC $0.27\pm 0.16$ $0.20\pm 0.07$ $0.22\pm 0.09$ PPO $0.91\pm 0.65$ $0.86\pm 0.43$ $0.76\pm 0.51$$\pm 10$ wind speed scenario RDDPG ${0.18\pm 0.04}$ ${0.18\pm 0.03}$ ${0.19\pm 0.05}$ DDPG $\mathbf{0.16\pm 0.06}$ $\mathbf{0.15\pm 0.07}$ $\mathbf{0.15\pm 0.08}$ TD3 $0.48\pm 0.39$ $0.38\pm 0.32$ $0.42\pm 0.33$ SAC $0.29\pm 0.16$ $0.25\pm 0.11$ $0.27\pm 0.13$ PPO $1.16\pm 0.82$ $1.13\pm 0.78$ $1.14\pm 0.77$$\pm 15$ wind speed scenario RDDPG $\mathbf{0.22\pm 0.07}$ $\mathbf{0.21\pm 0.07}$ $\mathbf{0.22\pm 0.07}$ DDPG $0.23\pm 0.15$ $0.23\pm 0.15$ $0.23\pm 0.15$ TD3 $0.56\pm 0.42$ $0.52\pm 0.37$ $0.55\pm 0.40$ SAC $0.46\pm 0.28$ $0.43\pm 0.26$ $0.43\pm 0.27$ PPO $1.50\pm 1.16$ $1.50\pm 1.12$ $1.50\pm 1.14$$\pm 20$ wind speed scenario RDDPG $\mathbf{0.22\pm 0.07}$ $\mathbf{0.21\pm 0.07}$ $\mathbf{0.22\pm 0.07}$ DDPG $0.37\pm 0.36$ $0.36\pm 0.36$ $0.35\pm 0.33$ TD3 $0.60\pm 0.47$ $0.57\pm 0.44$ $0.60\pm 0.46$ SAC $0.51\pm 0.36$ $0.51\pm 0.37$ $0.51\pm 0.38$ PPO $1.60\pm 1.28$ $1.57\pm 1.26$ $1.53\pm 1.20$
VII Conclusion

### v4

@@ 4.2 Robust deep deterministic policy gradient
    The corresponding target networks $\pi_{\theta^{\prime}}$, $\mu_{\phi^{\prime}}$ for the two online actor network
@@ 4.2.1 Critic update
$L_{\rm critic}(\psi_{i};B):=$ $\frac{1}{|B|}\sum\limits_{(x,u,w,c,x^{\prime})\in B}{(y-Q_{\psi_{i}}(x,u,w))}^{2},\quad i\in\{1,2\}$
@@
$\tilde{u}$ $=\pi_{\theta^{\prime}}(x^{\prime})+{\rm clip}({\epsilon_{1}},-\delta,\delta)\ \ \epsilon_{1}\sim\mathcal{N}(0,\sigma)$
$\tilde{w}$ $=\mu_{\phi^{\prime}}(x^{\prime})+{\rm clip}({\epsilon_{2}},-\delta,\delta)\ ,\ \epsilon_{2}\sim\mathcal{N}(0,\sigma)$
where $\epsilon_{1},\epsilon_{2}$ are random noise vectors added in order to smooth out the target values, $\mathcal{N}(0,\sigma)$ implies the Gaussian distribution with zero mean and variance $\sigma$, and the clip function is added in order to guarantee the boundedness of the control inputs. The critic’s online parameters $\psi_{1},\psi_{2}$ are updated by the gradient descent step to minimize the loss
${\psi_{i}}\leftarrow{\psi_{i}}-{\alpha_{{\rm{critic}}}}{\nabla_{\psi_{i}}}{L_{\rm critic}}({\psi_{i}};B),\quad i\in\{1,2\}$
@@ 4.2.2 Actor update
4.2.3 Target parameters update
@@ 4.2.4 Exploration
where $e_{1},e_{2}\sim\mathcal{N}(0,\sigma)$ are exploration noises [5]. The overall algorithm is summarized in Algorithm 1.
@@
[…]
8:   Select actions $u_{k}=\pi_{\theta}(x_{k})+e_{1}$ and $w_{k}=\mu_{\phi}(x_{k})+e_{2}$,
9:   where $e_{1},e_{2}\sim\mathcal{N}(0,\sigma)$ are exploration noises.
10:   Compute the cost $c_{k}:=c(x_{k},u_{k},w_{k})$
11:   Observe the next state $x_{k+1}$
12:   Store the transition tuple $(x_{k},u_{k},w_{k},c_{k},x_{k+1})$ in the replay buffer $D$
13:   Uniformly sample a mini-batch $B$ from the replay buffer $\mathcal{D}$
14:   Update critic network:
@@
15:   Update actor networks by the deterministic policy gradient:
@@
$\theta_{i}^{\prime}\leftarrow\tau\theta_{i}+(1-\tau)\theta_{i}$
$\phi_{i}^{\prime}\leftarrow\tau\theta_{i}+(1-\tau)\phi_{i}$
17:  end for
@@
5 Quadrotor application
5.1 Quadrotor dynamics model
@@
where $\mathbf{K}_{D}=\text{diag}(K_{D,x},K_{D,y},K_{D,z})$ is a matrix of drag coefficients. Among the four motors, two opposing motors ($1$th and 33rd) rotate clockwise, while the others (22nd and $4$th) rotate counterclockwise. This causes torque $\tau_{k}$ in the roll, pitch, yaw directions
$\tau_{k}=\begin{bmatrix}\tau_{k,\phi}\\ \tau_{k,\theta}\\ \tau_{k,\psi}\end{bmatrix}=\begin{bmatrix}l^{\prime}k_{F}(\text{RPM}_{k,2}^{2}-\text{RPM}_{k,4}^{2})\\ l^{\prime}k_{F}(-\text{RPM}_{k,1}^{2}+\text{RPM}_{k,3}^{2})\\ k_{M}(\text{RPM}_{k,1}^{2}-\text{RPM}_{k,2}^{2}+\text{RPM}_{k,3}^{2}-\text{RPM}_{k,4}^{2})\\ \end{bmatrix}$
where $l^{\prime}=\cos(\frac{\pi}{4})l$ is the scaled arm length, because the motors of the quadrotor are positioned at a 45-degree angle relative to the $x$ and $y$ axes when the yaw angle is 0, $k_{F}$ is the thrust coefficient, and $k_{M}$ is the torque coefficient. The inertia matrix, denoted by the symbol $I$, represents the inertial properties of the quadrotor’s body.
5.2 State space
@@
$x_{k}=[\mathbf{e}_{k}^{p},\mathbf{e}_{k}^{v},\boldsymbol{\rho}_{k},\boldsymbol{\omega}_{k}]^{T}\in\mathbb{R}^{12}$
where $\mathbf{e}_{k}^{p}$ denotes the position error, $\mathbf{e}_{k}^{v}$ represents the velocity error, $\boldsymbol{\rho}_{k}$ is the Euler angle vector, and $\boldsymbol{\omega}_{k}$ is the angular velocity vector. These quantities are defined as follows:
$\mathbf{e}_{k}^{p}$ $=\mathbf{p}_{k}^{\text{target}}-\mathbf{p}_{k}$
@@
$\mathbf{e}_{k}^{v}$ $=\mathbf{v}_{k}^{\text{target}}-\mathbf{v}_{k}$
@@
$\boldsymbol{\rho}_{k}$ $=[\phi_{k},\theta_{k},\psi_{k}]^{T}\in\mathbb{R}^{3}$
$\boldsymbol{\omega}_{k}$ $=[\omega_{k,\phi},\omega_{k,\theta},\omega_{k,\psi}]^{T}\in\mathbb{R}^{3}$
where $\mathbf{p}_{k}^{\text{target}}$ and $\mathbf{p}_{k}$ denote the position vectors of the target and the quadrotor, respectively, and $\mathbf{v}_{k}^{\text{target}}$ and $\mathbf{v}_{k}$ represent their corresponding linear velocity vectors. $\phi_{k}$$\theta_{k}$$\psi_{k}$ denotes roll, pitch, yaw of the quadrotor. As in the method used in DQN[2], we used the last four time steps of the state as input to the neural network to help the network better capture dynamics in the environment.
5.3 Action space
The action of the quadrotor, $u_{k}$, is the speed of four motors $\text{RPM}_{k,i},\ i\in\{1,2,3,4\}$, which is continuous value between 0 and $21713.714$, as follows:
@@ 5.4 Cost function
[…]
@@
where the scaling factors $\alpha$, $\beta$, $\epsilon$, and $\zeta$ are positive (i.e., $\alpha,\beta,\epsilon,\zeta>0$), and $\lambda$ is negative (i.e., $\lambda<0$). The specific values used for these parameters are summarized in the Appendix  2. Since $\alpha$ and $\beta$ are positive, the quadrotor agent learns a policy that minimizes position and velocity errors. If the quadrotor’s angles become too large, it becomes more susceptible to disturbances and may flip easily. Therefore, minimizing the angles cost $c_{\rho}$ is crucial for stability. Additionally, the living cost $c_{l}$ is a negative reward that encourages the quadrotor to maintain stability. The penalty cost $c_{tr}$ is a large positive value applied when the quadrotor fails to track the target—specifically, when the distance between the target and the quadrotor exceeds $5$ meters, or when the quadrotor flips with roll or pitch angles exceeding $\frac{\pi}{2}$. The agent learns a policy to avoid receiving this penalty, which can accelerate the initial learning process.
6 Experiments and results
6.1 Experiments setups
For experiments, we used the gym-pybullet-drones environment [17], which is an open-source quadrotor simulator built with Python and the Bullet Physics engine. This environment provided a modular and precise physics implementation, supporting both low-level and high-level control. In the simulator, the Crazyflie 2.0 quadrotor was modeled, and its parameters were summarized in Appendix Table 1.
To demonstrate the robustness of the proposed RDDPG algorithm, we compared it with other model-free reinforcement learning algorithms, including DDPG [3], TD3 [5], and SAC [6]. These methods utilized a single agent that aimed to minimize the cost function $\tilde{c}$. To enhance their robustness against disturbances, random external disturbances were applied during the training process. A detailed summary of the hyperparameters used in training was provided in Appendix 3 and  4.
@@ 6.1 Experiments setups
For evaluation, we generated two trajectories as shown in Fig. 2. During evaluation, we introduced random wind disturbances in the $x$ and $y$ directions. We considered three scenarios where the maximum wind velocity was limited to $\pm 5$, $\pm 10$, and $\pm 15$ [m/s], which were referred to as the $\pm 5$, $\pm 10$, and $\pm 15$ wind speed scenarios in the figures and tables. In each scenario, the initial wind speed was randomly set within a given range and updated every $10$ time steps with normally distributed noise.
Each algorithm was tested over 500 episodes and evaluated based on the mean and standard deviation of the total cost per episode, where the total cost was computed as the sum of the per-step costs $\tilde{c}(x_{k},u_{k})$ over a single episode. In addition, the mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor at each timestep were also considered as evaluation metrics.
Traj (1)
Traj (2)
Figure 2: Two evaluation trajectories, denoted as Traj (1) and Traj (2), shown from left to right.
6.2 Results
Figure 3: Mean and standard deviation of the total episode cost. RDDPG achieves the lowest average cost and the smallest variance across all scenarios. This indicates that the RDDPG yields a robust agent that performs consistently even under severe external disturbances.
[…]
By maintaining a consistently low and stable cost across a wide range of disturbance intensities, RDDPG exhibits superior robustness and reliable control performance, even under challenging environmental conditions. This enhanced robustness stems from its adversarial training framework, inspired by $H_{\infty}$ control theory and formulated as a two-player zero-sum game. By optimizing the policy against worst-case disturbances during training, RDDPG learns disturbance-resilient strategies that improve generalization and ensure stable performance under unseen or highly uncertain conditions.
Figure 4: Trajectory results of all algorithms under the $\pm 15$ wind scenario and Traj (1).
Figure 4 illustrates the trajectory tracking performance of each algorithm across varying wind speeds. RDDPG achieves the most accurate trajectory tracking, exhibiting the smallest tracking error compared to the other algorithms. The mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor is presented in Figure  5. Across all scenarios, RDDPG consistently achieves lower position errors compared to TD3 and SAC. In the case of DDPG, it exhibits lower position errors under mild wind disturbance scenarios. However, as the disturbance becomes stronger, its performance degrades significantly, resulting in higher errors compared to RDDPG. Furthermore, DDPG shows larger standard deviations, indicating less stable performance across episodes. Detailed numerical results are provided in at Appendix Table  5.
Figure 5: Mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor at each time step. RDDPG consistently achieves low error and standard deviation across all scenarios, whereas other methods exhibit unstable performance.
7 Conclusion